# Flow with FlorDB: Incremental Context Maintenance for the Machine Learning Lifecycle

### v1

@@ Abstract.
The metadata involved in integrating code, data, configuration, and feedback into predictive models is varied and complex. This complexity is further compounded by the agile development practices favored by data scientists and machine learning engineers. These practices emphasize high experimentation velocity and frequent deployments, which can make it challenging to keep track of all the relevant metadata Shankar et al. (2024); Garcia et al. (2018). The iterative nature of agile methods means that models, datasets, and configurations are constantly evolving. Each experiment might involve tweaks to the data preprocessing steps, changes in model hyperparameters, or updates to the deployment environment. The need for rapid iteration can lead to shortcuts or oversights in documentation and metadata management. Effective metadata management requires robust yet flexible tools and practices that can integrate and organize this information without slowing down the development process. Traditional context management often emphasizes a “metadata first” approach, which can introduce significant friction for developers. FlorDB reduces this friction through multiversion hindsight logging and incremental context maintenance, allowing developers to add and refine metadata after the fact Garcia et al. (2023). This “metadata later” approach enables a more flexible and incremental development process, allowing data scientists to focus on model creation and refinement without the burden of documentation upfront. As shown in a demo, FlorDB can be used to build AI/ML applications with integrated train-infer pipelines and managed feedback loops. Ultimately, the goal of FlorDB is to ensure that critical metadata is maintained accurately and efficiently, even in fast-paced agile workflows.
@@ 1.1. A Crisis of Metadata Management
One persistent challenge is balancing agility with the rigor of upfront documentation or “metadata first” approaches. While detailed documentation and metadata are essential for reproducibility, debugging, and collaboration, insisting on comprehensive metadata from the start can hinder the agility and speed that are often crucial in early development stages. This balance is critical in enabling teams to innovate rapidly while ensuring long-term project maintainability and scalability.
At the heart of this crisis lies an enduring open problem: the tradeoff between discipline and agility. This dilemma is evident in various contexts, such as the dichotomy between schema-first and NoSQL databases, and the contrast between data warehouses and lakehouses. Schema-first approaches offer discipline and structure but lack the flexibility needed for rapidly evolving requirements. Conversely, NoSQL databases provide the necessary agility but can lead to inconsistency and disarray in metadata management.
To address this, two core goals can be established:
@@
    Goal 1. Agile Development Loop: Metadata capture should be gradually incorporated into the agile workflows of data scientists and MLEs without interference. For data scientists, it should fit naturally into their open-source development environment, including tools like Python, Jupyter, and Git. For MLEs, it should leverage standard tools for workflow management and scalability, such as relational databases and CI/CD pipelines, without requiring lock-in on yet another service for metadata management.
@@
    Goal 2. Metadata on Demand: Metadata generation, like other features, should be able to evolve incrementally, improving organically with project needs. Ideally we want a “metadata later” approach that easily back-propagates metadata discipline into earlier versions of the project — at just the time developers regret not having it.
@@ 1.2. Goals and Contributions
The scope of FlorDB [^1] https://github.com/ucbrise/flor has been significantly expanded to encompass the entire ML lifecycle, covering complete pipelines and their regular execution, rather than just individual tasks. These extensions apply to mechanisms managed by the FlorDB API, initially designed for multiversion hindsight logging but now adapted to include incremental context maintenance over workflows. Despite these extensions, FlorDB maintains API stability and backward compatibility with multiversion hindsight logging.
@@
    Incremental Context Maintenance for Agile DevOps: Our system allows developers to log and analyze metadata in a standard, open, low-friction manner. Metadata can be captured naturally through log statements as part of the development workflow, without imposing significant overhead. Subsequently, these log statements can be read back directly as tabular data using a standard Python dataframe library, Pandas, without any need for data wrangling.
@@
    Multiversion Hindsight Logging for “Metadata on Demand”: Backward-compatibility maintains FlorDB support for multiversion hindsight logging Garcia et al. (2023). Our multiversion hindsight logging mechanism eliminates the need for “metadata first”. Developers can add or refine metadata post-hoc, on demand. This “metadata later” approach ensures that developers can focus on development and rapid iteration while still maintaining the benefits of structured metadata management.
Last, the open, standard approach of FlorDB simplifies and improves the abstraction of metadata, incorporating features from various bespoke ML metadata systems such as feature stores, model registries, and labeling systems into a unified and robust framework. We demonstrate this through a document intelligence use-case, highlighting the importance of context in streamlining development cycles and improving operational efficiency.
@@ 2. Multiversion Hindsight Logging
This section provides background on Flor and FlorDB, highlighting their evolution and key features in managing the machine learning lifecycle. Flor, originally designed as a record-replay system for model training Garcia et al. (2021), offers two main features: i) low-overhead adaptive checkpointing, minimizing computational resources during model training, and ii) low-latency replay from checkpoints, leveraging memoization and parallelism speed ups. Flor record-replay added support for hindsight logging, allowing for post-hoc and on-demand querying of effectively unbounded context.
@@
FlorDB extends Flor’s capabilities by integrating automatic version control, adding a relational data model for querying logs, and cross-version logging statement propagation for multiversion hindsight logging Garcia et al. (2023). Its relational model, accessible via flor.dataframe, maps individual logging statements into columns in a pivoted view. This approach facilitates easy tracking of changes over time. In this paper, we further extend FlorDB to support dataflow pipelines and manage feedback loops. This extension provides a seamless framework for defining and executing complex, evolving ML pipelines.
2.1. Metadata on Demand
One of the key advantages of Multiversion Hindsight Logging is its support for metadata on demand, offering significant flexibility in capturing and utilizing metadata. This approach allows for the generation and refinement of metadata as needed, rather than necessitating comprehensive documentation upfront. This flexibility is especially valuable in agile environments, where rapid iteration and frequent changes are common.
In traditional systems, metadata must often be defined and captured early in the project lifecycle. However, in practice, data scientists and MLEs frequently refine their processes, models, and tools based on new insights and evolving project requirements. Multiversion Hindsight Logging accommodates these changes by enabling the capture of metadata at any stage of the project’s lifecycle. The ability to log and query metadata retrospectively means that developers can revisit and refine metadata after the initial development phases. This is particularly useful for capturing metadata that may have been overlooked or considered unnecessary in the early stages. As the project matures and the need for detailed documentation increases, developers can retroactively add and refine metadata.
FlorDB’s approach to metadata on demand provides a flexible and efficient solution to the challenges of metadata management in agile ML development. It balances the need for rapid iteration with the benefits of thorough documentation and metadata capture, ensuring that critical metadata is both available when needed and manageable over time. The changes and extensions presented in this paper are fully backward-compatible with FlorDB’s multiversion hindsight logging capabilities, extending metadata on demand to entire ML pipelines, not just individual training scripts.
@@ 2.2. FlorDB Extended API
FlorDB’s API captures metadata about the executing file, eliminating the need to restate dataflow dependencies; the Makefile suffices. By profiling runtime metadata, including the executed file’s name, FlorDB remains agnostic to the workflow or dataflow management system, functioning seamlessly whether called by Make or Airflow (or others) without requiring refactoring.
@@
    A context manager defining objects for adaptive checkpointing at flor.loop iteration boundaries.
@@ 3. Incremental Context Maintenance
Our conceptualization of context is drawn from the work of Hellerstein et al. (2017) Hellerstein et al. (2017), who proposed it as an extension of traditional database metadata. They introduced the “ABCs of Context” mnemonic, which we find particularly useful for understanding and navigating the complex landscape of ML metadata. This framework provides a structured approach to capturing and managing the rich tapestry of information that underpins real-world ML applications.
3.1. Contextualized Logging
“Application Context” represents core information describing what raw data an application processes and interprets Hellerstein et al. (2017). This encompasses all information that could be logged, such as the values of arbitrary expressions at runtime, which FlorDB can capture post-hoc with multiversion hindsight logging (Section 2) and manages using a unified API, providing a system that supports flexible, NoSQL-like data writes and powerful, SQL-like data reads.
FlorDB provides a straightforward interface for logging via flor.log(name, value) statements, ensuring that each log entry is accompanied by crucial metadata such as projid, tstamp, filename, and ctx_id. This metadata is captured at the time of import and embedded within every log entry, enabling unambiguous identification of the log’s origin and context. The ctx_id is generated during the initialization of flor.loop and indicates the specific loop context a log entry belongs to, providing visibility into nested operations and possible cross-iteration dependencies (see logs and loops in Figure 1).
In Figure 2, we give an example of how FlorDB’s logging mechanism captures data features during a document analysis process. The example illustrates how multiple documents are processed, each consisting of several pages. This example shows how FlorDB efficiently captures a wide range of metadata without requiring a predefined schema. The resulting data, including the logs of headings, page numbers, text sources, and page texts, is then accessible through the flor.dataframe. As shown in the bottom part of Figure 2, the dataframe presents this metadata in a structured layout, allowing users to query and analyze the data more efficiently Harinarayan et al. (1996).
1for doc_name in flor.loop("document", os.listdir(...)):
2 N = get_num_pages(doc_name)
3 for page in flor.loop("page", range(N)):
4 # text_src is "OCR" or "TXT"
5 text_src, page_text = read_page(doc_name, page)
6 flor.log("text_src", text_src)
7 flor.log("page_text", page_text)
8
9 # Run some featurization
10 headings, page_numbers = analyze_text(page_text)
11 flor.log("headings", headings)
12 flor.log("page_numbers", page_numbers)
Figure 2. Data featurization with FlorDB

### v2

@@ Abstract.
In this paper we present techniques to incrementally harvest and query arbitrary metadata from machine learning pipelines, without disrupting agile practices. We center our approach on the developer-favored technique for generating metadata — log statements — leveraging the fact that logging creates context. We show how hindsight logging Garcia et al. (2021) allows such statements to be added and executed post-hoc, without requiring developer foresight. Relational views of incomplete metadata can be queried to dynamically materialize new metadata in bulk and on demand across multiple versions of workflows. This is done in a “metadata later” style, off the critical path of agile development. We realize these ideas in a system called FlorDB and demonstrate how the data context framework covers a range of both ad-hoc metadata as well as special cases treated today by bespoke feature stores and model repositories. Through a usage scenario—including both ML and human feedback—we illustrate how the component techniques come together to resolve classic software engineering trade-offs between agility and discipline.
@@ 1.1. A Crisis of Metadata Management
One persistent challenge is balancing agility with the rigor of upfront documentation or “metadata first” approaches. Detailed documentation and metadata are essential for reproducibility, debugging, and collaboration. Unfortunately, insisting on comprehensive metadata from the start can hinder the agility and speed that are often crucial in early development stages.
The tradeoff between discipline and agility is an enduring open problem, evident in various contexts including the dichotomy between schema-first and NoSQL databases, and the contrast between data warehouses and data lakes. Schema-first and warehouse approaches offer discipline and structure but lack the flexibility needed for rapidly evolving requirements. Conversely, NoSQL and data lake approaches provide the necessary agility but can lead to inconsistency and disarray in metadata management.
To address this, two core goals can be established for resolving what has been considered an inherent conflict:
@@
    Goal 1. Agile Development Loop: Metadata capture should be gradually incorporated into the agile workflows of data scientists and MLEs without interference. For data scientists, metadata capture should fit naturally into their open-source development environment. For MLEs, metadata should fit into standard tools for workflow management and scalability, such as relational databases and CI/CD pipelines, without requiring lock-in on yet another service for metadata management.
@@
    Goal 2. Metadata on Demand: Metadata generation, like other features, should be able to evolve incrementally, improving organically with project needs. Ideally we want a “metadata later” approach that easily back-propagates metadata discipline into earlier versions of the project — at just the time when someone regrets not having the metadata they need.
@@ 1.2. Goals and Contributions
Our work is built on FlorDB [^1] https://github.com/ucbrise/flor, the scope of which we significantly expanded to encompass the entire ML lifecycle. FlorDB now captures complete pipelines and their regular execution, rather than just individual tasks. These extensions apply to mechanisms managed by the FlorDB API, initially designed for multiversion hindsight logging but now adapted to include incremental context maintenance over workflows. Despite these extensions, FlorDB maintains API stability and backward compatibility with multiversion hindsight logging.
@@
    Incremental Context Maintenance for Agile DevOps: Our system allows developers to log and analyze metadata in a standard, open, low-friction manner. Metadata can be captured naturally through Python log statements as part of the development workflow, without imposing significant overhead. Subsequently, these log statements can be read directly as tabular data using standard Python dataframes, queried via Pandas or SQL, without requiring data wrangling.
@@
    Ad-Hoc Metadata on Demand Enabled by Hindsight Logging: FlorDB enables agile metadata evolution through multiversion hindsight logging, which operates as a record-replay mechanism. While traditional logging systems require predefined schemas, our approach captures sufficient execution state during recording, and allows the extraction of arbitrary expression values derivable from that state during replay. This means metadata collection isn’t limited to just what was explicitly recorded; developers can compute arbitrary new properties and relationships from the execution record after the fact. Like materializing a view, developers can declare what metadata they want to extract, but unlike views, the source material extends beyond stored data to encompass both the complete execution record and the universe of derivable properties. This powerful mechanism frees developers from “metadata first” constraints while enabling unbounded possibilities for post-hoc metadata computation and analysis.
  3. (3)
    Unified, Open Metadata for Machine Learning: Last, the open, standard approach of FlorDB simplifies and improves the abstraction of metadata, incorporating features from various bespoke ML metadata systems such as feature stores, model registries, and labeling systems into a unified and robust framework.
We demonstrate these contributions through a document intelligence use-case, highlighting the importance of context in streamlining development cycles and improving operational efficiency.
@@ 2. Multiversion Hindsight Logging
This section provides background on Flor and FlorDB, highlighting their evolution and key features in managing the machine learning lifecycle. Flor, originally designed as a record-replay system for model training Garcia et al. (2021), offers two main features: i) low-overhead adaptive checkpointing, minimizing computational resources during model training, and ii) low-latency replay from checkpoints, leveraging memoization and parallelism speed ups. Flor’s record-replay mechanism introduced the notion of hindsight logging, allowing for post-hoc and on-demand querying of effectively unbounded context.
We adopt the term “context” from the Ground project Hellerstein et al. (2017), as a signifier for an all-encompassing view of metadata that goes beyond traditional relational metadata. Here, “context” includes anything that could be emitted by a log statement in any running process, be it an ML training job, a data wrangling script, a live inference server, a log processor handling usage feedback, or even an orchestration framework that knits these other tools together.
@@
FlorDB extends Flor’s capabilities by integrating automatic version control, adding a relational data model for querying logs, and cross-version logging statement propagation for multiversion hindsight logging Garcia et al. (2023). Its relational model, accessible via flor.dataframe, maps individual logging statements into columns in a pivoted view. This approach facilitates easy tracking of changes over time. In this paper (Section 2.1), we further extend FlorDB to support dataflow pipelines and manage feedback loops. This extension provides a seamless framework for capturing arbitrary context introduced in defining and executing complex, evolving ML pipelines.
To clarify the concept of multiversion hindsight logging, imagine a scenario where a developer has run several versions of a machine learning pipeline but later realizes that certain metadata or context was not captured during those runs. Traditionally, retrieving this missing information would require re-running each version of the pipeline with the new logging statements inserted—a process that is both time-consuming and resource-intensive. FlorDB’s multiversion hindsight logging offers a powerful alternative that minimizes both developer effort and compute resources. Developers can add the desired logging statements to the latest version of their code, and FlorDB will (a) inject these statements into the correct locations in all prior versions of the code, and (b) retroactively execute these statements across all those versions via incremental replay, without the need for full re-execution. The former, (a), is made possible via techniques adapted from code diffing Falleri et al. (2014); the latter, (b), is made possible through a combination of differential execution and parallelism, allowing FlorDB to efficiently replay only the necessary parts of the pipeline to extract the new metadata. This “magic trick” enables developers to incrementally build up context and metadata after the fact, supporting agile practices by eliminating the need for foresight in logging. It resolves the classic trade-off between starting fast and refining over time by providing the flexibility to log now, and get data from the past. For technical details on how this is achieved, we refer readers to our prior work Garcia et al. (2023).
@@ 2.1. FlorDB Extended API
FlorDB’s API captures metadata about the executing file, eliminating the need to restate dataflow dependencies; cross-executable dependencies declared in a typical workflow orchestration tool (Airflow, MLFlow, Make, etc) suffice. By profiling runtime metadata, including the executed file’s name, FlorDB remains agnostic to the choice of workflow management system, functioning seamlessly without requiring refactoring of orchestration scripts.
@@
    A Python "context manager" defining objects for adaptive checkpointing at flor.loop iteration boundaries.
@@ 3. Incremental Context Maintenance
We base our conceptualization of context on the “ABCs of Context” framework introduced by Hellerstein et al. (2017) Hellerstein et al. (2017), which extends traditional database metadata to encompass a broader spectrum of information critical in ML applications. The framework includes:
  1. (1)
    Application Context (the “A”): Captures how raw data is interpreted, including schemas, checkpoints, and parameters. FlorDB captures application context through log statements, allowing developers to record pertinent information during execution.
  2. (2)
    Behavioral Context (the “B”): Tracks how data is created and used over time, relating to lineage or provenance. This context is often contained in build files, capturing dependencies and directed acyclic graphs (DAGs) of tasks. FlorDB captures behavioral context via its relational data model, associating each logging statement with its originating filename. This mapping reveals dataflow pathways through the codebase, enabling analysis of data transformations and lineage within the pipeline.
  3. (3)
    Change Context (the “C”): Manages version histories of both data and code. FlorDB manages change context using Git version control, ensuring that all runs and modifications are tracked and retrievable.
This framework provides a structured approach to understanding and managing the rich tapestry of information that underpins real-world ML applications.
@@
3.1. Application Context
Application Context represents core information describing what raw data an application processes and interprets Hellerstein et al. (2017). This includes all information that could be logged, such as the values of arbitrary expressions at runtime. FlorDB can capture this information post-hoc using multiversion hindsight logging (Section 2) and manages it through a unified API. This approach provides a system that supports flexible, NoSQL-like data writes and powerful, SQL-like data reads.
FlorDB provides a straightforward interface for logging via flor.log(name, value) statements, ensuring that each log entry is accompanied by crucial structured metadata such as projid, tstamp, filename, and ctx_id. This metadata is captured at the time of import and embedded within every log entry, enabling unambiguous identification of the log’s origin and context. The ctx_id is generated during the initialization of flor.loop and indicates the specific loop context a log entry belongs to, providing visibility into nested operations and possible cross-iteration dependencies (see logs and loops in Figure 1).
In Figure 3, we give an example of how FlorDB’s logging mechanism captures data features during a document analysis process. The example illustrates how multiple documents are processed, each consisting of several pages. This example shows how FlorDB captures a wide range of metadata without requiring a predefined schema. The resulting data, including the logs of headings, page numbers, text sources, and page texts, is then accessible through the flor.dataframe. As shown in the bottom part of Figure 3, the dataframe presents this metadata in a structured layout, allowing users to query and analyze the data more efficiently Harinarayan et al. (1996).
1for doc_name in flor.loop("document", os.listdir(...)):
2 N = get_num_pages(doc_name)
3 for page in flor.loop("page", range(N)):
4 # text_src is "OCR" or "TXT"
5 text_src, page_text = read_page(doc_name, page)
6 flor.log("text_src", text_src)
7 flor.log("page_text", page_text)
8
9 # Run some featurization
10 headings, page_numbers = analyze_text(page_text)
11 flor.log("headings", headings)
12 flor.log("page_numbers", page_numbers)
Figure 3. Data featurization with FlorDB
3.2. Behavioral & Change Context
Behavioral and change context are naturally intertwined in ML pipelines: every execution both advances through the dependency graph (behavioral) and creates a new version in time (change). This context emerges naturally in FlorDB via logging executions with relevant file names and timestamps, rather than requiring manual documentation effort integrated into a workflow manager.
Behavioral context captures how data is created and used: dependency management, provenance and lineage, pipeline pathways, and dataflow. Change context tracks the version history of data, code, configuration parameters, checkpoints, and associated information Hellerstein et al. (2017). Together, they help answer critical development questions: Where was this data defined? Where is it transformed? Who made the last change? Where should new transformations be added?
Currently, teams often manage these contexts through face-to-face interactions or communication tools like Slack or email Shankar et al. (2024). However, this approach creates friction by relying on colleagues’ availability and memory, and doesn’t scale as teams change or projects evolve. While some teams try to maintain comprehensive documentation, this conflicts with the rapid iteration typical in MLOps environments.
FlorDB manages both contexts through its logging system. When users query a flor.dataframe, they receive not just data but its lineage and version history. To illustrate how FlorDB handles behavioral and change context, let’s consider a simplified version of our document intelligence pipeline (we will cover this case in more depth in Section 4).
process_pdfs: \$(PDFS) pdf_demux.py
@echo "Processing PDF files..."
@python pdf_demux.py
@touch process_pdfs
featurize: process_pdfs featurize.py
@echo "Featurizing Data..."
@python featurize.py
@touch featurize
train: featurize hand_label train.py
@echo "Training..."
@python train.py
model.pth: train export_ckpt.py
@echo "Generating model..."
@python export_ckpt.py
infer: model.pth infer.py
@echo "Inferencing..."
@python infer.py
@touch infer
hand_label: label_by_hand.py
@echo "Labeling by hand"
@python label_by_hand.py
@touch hand_label
run: featurize infer
@echo "Starting Flask..."
@flask run
Our simplified pipeline comprises three main Python scripts that interact with FlorDB:
  1. (1)
    train.py: This script fine-tunes a machine learning model using the preprocessed data. It loads the dataset prepared by prep.py, trains the model, and saves a model checkpoint along with performance metrics like accuracy and recall. Throughout the process, it logs the training data, model parameters, and metrics, capturing the change context over time.
  2. (2)
    infer.py: This script performs inference using the most effective model checkpoint. flor.dataframe("acc", "recall") is queried to retrieve the model checkpoint with the highest recall from the execution history. By accessing the logged performance metrics and version timestamps, it ensures that predictions are made using the best available model. The script then processes new data and logs the predictions for further analysis.
  3. (3)
    run.py: This script launches a Flask web application that serves the model’s predictions to end-users and collects human feedback. Users can review and correct the model’s outputs, providing valuable annotations. The script logs these interactions using flor.log(name, value), capturing both the behavioral context (how data flows through the system) and the change context (how feedback updates the dataset).
The dependencies between these scripts are specified in a Makefile, which orchestrates the execution order based on these dependencies (left pane, Figure 2). This dependency management ensures that each script runs in the correct order and that each stage has the necessary data and models from the previous stages. While we’ve used a Makefile for simplicity, these dependencies can also be specified using other workflow management tools like Airflow or MLflow, depending on the project’s complexity and requirements.
In this pipeline, the process cycles between collecting new human-reviewed data with run.py and updating the model with train.py. The infer.py script ensures that the inference stage always utilizes the best-performing model checkpoint.
Figure 2 illustrates the flow of data and transformations in the pipeline, highlighting how FlorDB captures both behavioral and change contexts. By logging with FlorDB, machine learning engineers (MLEs) not only build pipelines but also maintain a comprehensive history of data transformations and model updates. This approach reduces reliance on ad-hoc communication and documentation, allowing projects to evolve without losing crucial context.
4. PDF Parser Demo
Real-world AI/ML applications are backed by ML pipelines of non-trivial complexity, often encompassing both computation and human-in-the-loop feedback. This section gives an overview of the PDF Parser Demo, a practical application of FlorDB in document intelligence. We demonstrate how FlorDB can be effectively used to manage context and dataflow that spans multiple asynchronous tasks that can generate a variety of metadata via computation and human feedback.
The PDF Parser [^2] https://github.com/ucbepic/pdf_parser is a Flask-based web application designed for efficient PDF document processing, including tasks like splitting PDFs, extracting text, and preparing data for analysis using natural language processing (NLP) techniques. Users interact with the parser through a simple web interface, as shown in Figure 4 (left). This interface allows users to navigate PDF documents and select the desired processing options.
In this demo, we aim to achieve several goals that showcase the capabilities of FlorDB:
  • Demonstrate Basic Functionality:
    • Version Tracking Across Code Changes: We will show how FlorDB’s multiversion hindsight logging operates seamlessly even when the code has been refactored, ensuring that all changes are reviewable.
    • Flexible Pipeline Modification: The demo illustrates how new stages can be inserted into an existing pipeline or new pathways can be incorporated into a directed acyclic graph (DAG) within the Flor data model, enabling dynamic workflow evolution.
    • Incremental Replay Execution: By using the Flor dataframe, we can drive incremental replay execution, re-running only the parts of the workflow that have been selected by the user, thus saving time and resources.
  • Showcase FlorDB Across Multiple Contexts:
    • Feature Storage and Querying After Execution: We demonstrate how FlorDB can store and allow queries on data features post-execution without prior setup.
    • Model Registry Functionality Post-Execution:
        Comparative Metrics Across Models: The demo compares performance metrics across different training runs after they have been executed.
        Post-Hoc Governance Enforcement: We show how to apply governance policies retroactively to identify and handle issues like corrupted or malicious datasets (e.g., detecting a poisoned dataset).
    • Metric Registry and Visualization After Execution: FlorDB acts as a repository for metrics post-execution, allowing for visualization of results—similar to generating and augmenting TensorBoard plots—even if this wasn’t configured beforehand.
Beyond highlighting the practical value of a context-rich approach, we hope this demo will serve as a reference implementation for those looking to get started with FlorDB.
4.1. PDF Extraction & Text Featurization
Once the PDF is converted into text and image formats, ensuring there is one document per page, the process of featurization begins (Figure 3). This process typically involves text extraction, feature engineering, and vectorization. This featurization process is essential for transforming raw PDF data into a structured form that is amenable to analysis and machine learning applications. The described methodology focuses on maximizing the information extracted from each page, ensuring that both textual and visual data contribute to the inferences made on the document. Takeaway: When used in featurization contexts (Figure 3), FlorDB can provide the functionality of a feature store.
4.2. Inference Pipeline
The inference pipeline automates the processing and analysis of images organized into document-specific folders. FlorDB’s features enhance this pipeline in several key ways. First, FlorDB’s comprehensive versioning and metadata tracking enables intelligent model selection. Through queries to FlorDB’s unified metadata store (e.g., flor.dataframe("acc", "recall")), the pipeline can automatically select the best-performing model checkpoint based on validation metrics tracked across all training runs. This eliminates the need for manual record-keeping or separate model registries.
Second, FlorDB’s logging infrastructure streamlines the processing of document pages and images. As each image is processed, pre-processed, and analyzed, FlorDB captures the full context of transformations and model predictions. This includes tracking input parameters, preprocessing steps, and model outputs in a way that maintains clear lineage between the original documents and their derived predictions.
Third, when issues arise in production, FlorDB’s hindsight logging capability allows developers to retroactively add logging statements to debug problems without re-running the entire pipeline. The model then makes predictions on the images, with FlorDB logging the final inferences alongside their complete provenance. Takeaway: When used in inference pipelines, FlorDB functions as a comprehensive model registry (dataframe in Figure 2), managing not only checkpoint selection but also model versioning, metadata tracking, and performance metrics. This enables robust model lifecycle management while maintaining clear provenance of how models evolve through training iterations.
4.3. Training Pipeline
The training pipeline encapsulates a typical machine learning workflow, tailored for classifying images extracted from PDF pages (Figure 5). This pipeline performs a load of training data (line 1 in Figure 5, managed by FlorDB), and data preparation. A model architecture is defined, and the model is trained over a number of epochs. Performance is monitored by flor logging, and context is managed and tracked using FlorDB. Takeaway: When used in training pipelines, FlorDB can function as a training data store (line 1 in Figure 5), and a model repository (lines 3-21 in Figure 5).